In a RC circuit of R = 1000 Ωand C= 500 μF how long does it take to charge the capacitor to 50% of its maximum?
Answer #1

Time Constant is given by

T=RC=1000*(500*10-6)=0.5 s

In a RC circuit charge as a function of time for charging is given by

Q=Qmax(1-e-t/T)

0.5Qmax = Qmax(1-e-t/0.5)

Ln(0.5)=-t/0.5

t=0.3466 s